OpenStudy (anonymous):

Solve Each inequality. Graph the solution set and write the answer in interval notation. I am Terrible at math please walk me through the steps 4 less than or equal to k+11 over 4 less then or equal to 4

OpenStudy (anonymous):

\[4\le \frac{ k+11 }{ 4 }\le4\]

OpenStudy (anonymous):

\[4\times4\le4\times \frac{ k+11 }{ 4}\le4\times4\]

OpenStudy (anonymous):

\[16\le11K \le16\]